How many gallons of pure water should be evaporated from 160 gallons of a solution containing 12% salt so that the remaining solution contains 20% salt?
Answer by ptaylor(2198)

How many gallons of pure water should be evaporated from 160 gallons of a solution containing 12% salt so that the remaining solution contains 20% salt?
Let x=amount of pure water that needs to be evaporated
Now we know that the amount of salt that exists before the evaporation takes place (0.12)(160) has to equal the amount of salt that exists after the evaporation takes place (0.20)(160-x). Sooooo
(0.12)(160)=(0.20)(160-x) simplify
19.2=32-0.20x
0.20x=12.8
x=64 gal----amount of pure water that needs to be evaporated
CK
(0.12)(160)=(0.20)(96)
19.2=19.2
Hope this helps--ptaylor